SOLVE THE FOLLOWING SYSTEM (USE THE METHOD YOU PREFER, WHETHER IT BE ADDITION OR SUBSTITUTION)
6) y-3x= -2
6x-2y= 4
on this one i think i should use substitution, altough i'm really confused on this process.. please help!!
thanks,
lj23kfuller@yahoo.com

Answer by stanbon(57347)

y-3x= -2
6x-2y= 4
---------------
Substitute to get:
6x -2*(3x-2) = -2
6x-6x+4 = -2
4 = -2
---------------
This is a contradiction that resulted from assuming
the x's and the y's in the two equations were the same for
both equations.
----------------
That assumption lead to the contradiction.
Therefore the x's and the y's cannot be the same
in the two equations. That means there is no solution
for the system.
------------------
That also means the system of equations is "inconsistent".
---------------------
If you graphed the two equation you would get parallel lines.
=====================
Cheers,
Stan H.